# Physics-informed Neural Mapping and Motion Planning in Unknown Environments

### v1

@@ V-B Mapping Comparison
Figure 5: Comparison for motion planning in Gibson environments. The figures show five paths generated by our method (orange), LazyPRM (gray), iSDF+MPOT (red), KinectFusion+MPOT (cyan), iSDF+RRTConnect (green), KinectFusion+RRTConnect (yellow), nvblox+MPOT (blue), and nvblox+RRTConnect (pink). The table shows statistical results on 8$\times$100 different starts and goals in eight Gibson environments. We load grid-based maps in both CPU and GPU in RRTConnect to test the performance.
@@ V-C Motion Planning Comparison
We randomly choose 100 start and goal points in each environment and run the above-mentioned motion planning methods to get the final paths. We compare the path length, computational time, and success rate. The collision checking is performed using the ground-truth map to validate the generated paths and compute the success rate. We also set a time limit of 10 seconds for sampling-based methods LazyPRM and RRTConnect, after which the case is considered a failure. Fig. 5 shows the paths where our method and KinectFusion+MPOT generate short and smooth paths, but iSDF+MPOT and nvblox+MPOT generate damping results. RRTConnect only considers the SDF zero-level-set; thus, all RRTConnect generate similar results. Table II presents the statistical results. Our method’s computational time is about 30 times faster than MPOT and achieves 98% success rate. In addition, our method generates smooth paths with a safe margin to obstacles, while RRTConnect generates paths close to the obstacle, resulting in relatively shorter lengths. Additionally, iSDF is slower due to its slower collision query through the neural network compared to a grid-based collision query. In summary, our results validate that our online reconstructed Active NTFields achieve the best performance for motion planning.

### v2

@@ V-B Mapping Comparison
Figure 5: Comparison for motion planning in Gibson environments. The figures show five paths generated by our method (orange), LazyPRM (gray), iSDF+MPOT (red), KinectFusion+MPOT (cyan), iSDF+RRTConnect (green), KinectFusion+RRTConnect (yellow), nvblox+MPOT (blue), and nvblox+RRTConnect (pink). The table shows statistical results on 8$\times$100 different starts and goals in eight Gibson environments. We load grid-based maps in both CPU and GPU in RRTConnect to test the performance. The highlighted region illustrates that MPOT and RRTConnect may discover longer paths belonging to a different homotopy class.
@@ V-C Motion Planning Comparison
We randomly choose 100 start and goal points in each environment and run the above-mentioned motion planning methods to get the final paths. We compare the path length, computational time, and success rate. The collision checking is performed using the ground-truth map to validate the generated paths and compute the success rate. We also set a time limit of 10 seconds for sampling-based methods LazyPRM and RRTConnect, after which the case is considered a failure.
Fig. 5 shows the paths where our method and KinectFusion+MPOT generate short and smooth paths, but iSDF+MPOT and nvblox+MPOT generate damping results. RRTConnect only considers the SDF zero-level-set; thus, all RRTConnect generate similar results. Table II presents the statistical results. Our method’s computational time is about 30 times faster than MPOT and achieves 98% success rate. In addition, our method generates smooth paths with a safe margin to obstacles, while RRTConnect generates paths close to the obstacle, resulting in relatively shorter lengths.
All sampling-based methods failed in some cases primarily due to exceeding the time limit. These methods rely on sampling strategies that can require significant time to explore complex environments and find a global collision-free path, especially when narrow passages are present. In contrast, MPOT failed primarily because it was unable to find a global collision-free path, even when sufficient time was available, as it depends on local optimization, which may get trapped in local minima without a global guarantee.
Our method, like MPOT, also failed in some cases because it could not find a global collision-free path, but it demonstrated higher robustness and faster convergence. To clarify the success rate further, failures for sampling-based methods occurred due to exceeding the time limit, while for MPOT and our method, failures were due to the inability to find a feasible path without collisions.